1 how do you solve 3/7 of what number is 9?
2 how do you solve 2/3 of what number is 16
3 solve for z 6=z/5.9
4what is trhe nearest tenth for 449/0.42

Answer by rfer(16322)

9/(3/7)=X
9*7/3=63/3
x=21
---------------
16/(2/3)=x
16*3/2=x
x=24
--------------
6=z/5.9
6*5.9=z
z=35.4
-------------
1069.0